OpenStudy (anonymous):

Another process of elimination question! State if there are no solutions or if there are infinite solutions. 8x + 10y = 92 24x + 11y = 162

OpenStudy (anonymous):

8x + 10y = 92 Multiply both sides by 3 24x+30y=276 24x+30y=276 24x + 11y = 162 19y=276-162=114=19*6 y=6 Plug y=6 into one of the equations to get x.

OpenStudy (anonymous):

thank you so much! I plugged it in, got x =4 and then checked it! I had forgotten originally to multiply by 3! Thanks now I remember the format for next time!

OpenStudy (anonymous):

You're welcome^^